A 9-year zero coupon bond with a $1,000 face value has an interest rate of 3.6% per year. What would be the change in the bonds value if the 9-year interest rate were to rise by 34 basis points. (Remember: your answer should not quote in percent or basis points.)
Answer #1

Price of the bond now=1000/1.036^9=727.3814165

Price of the bond after rate increase=1000/(1+3.6%+0.34%)^9=706.2453276


Change=706.2453276-727.3814165=-21.1360889

% change=-21.1360889/727.3814165=-2.90578% or -290.578 basis points